# ViSurf: Visual Supervised-and-Reinforcement Fine-Tuning for Large Vision-and-Language Models

### v2

@@ Abstract
Typical post-training paradigms for Large Vision-and-Language Models (LVLMs) include Supervised Fine-Tuning (SFT) and Reinforcement Learning with Verifiable Rewards (RLVR). SFT leverages external guidance to inject new knowledge, whereas RLVR utilizes internal reinforcement to enhance reasoning capabilities and overall performance. However, our analysis reveals that SFT often leads to sub-optimal performance, while RLVR struggles with tasks that exceed the model’s internal knowledge base. To address these limitations, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified post-training paradigm that integrates the strengths of both SFT and RLVR within a single stage. We analyze the derivation of the SFT and RLVR objectives to establish the ViSurf objective, providing a unified perspective on these two paradigms. The core of ViSurf involves injecting ground-truth labels into the RLVR rollouts, thereby providing simultaneous external supervision and internal reinforcement. Furthermore, we introduce three novel reward control strategies to stabilize and optimize the training process. Extensive experiments across several diverse benchmarks demonstrate the effectiveness of ViSurf, outperforming both SFT, RLVR, and two-stage SFT → RLVR. In-depth analysis corroborates these findings, validating the derivation and design principles of ViSurf. The code will be publicly available.
@@ 1 Introduction
Developing Large Vision-and-Language Models (LVLMs) that excel in diverse visual perception tasks is a promising direction for visual intelligence. To this end, prior work has predominantly relied on two training paradigms: Supervised Fine-Tuning (SFT) [liu2023visual, wang2024qwen2vl, bai2025qwen25vl] and Reinforcement Learning with Verifiable Rewards (RLVR) [liu2025seg, liu2025visionreasoner].
@@
Every coin has two sides, and these paradigms are no exception. SFT directly optimizes the model using expert-annotated data. This approach provides explicit external guidance, enabling the model to memorize the target distribution. However, it often exhibits sub-optimal performance and can lead to catastrophic forgetting of pre-trained knowledge. Recently, RLVR has garnered significant research attention. The on-policy RLVR methods start with an initial policy and iteratively generates a group of rollouts that are evaluated by pre-defined reward functions, then the policy is subsequently optimized based on this internal feedback, such as Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] and Dynamic Sampling Policy Optimization (DAPO) [yu2025dapo]. By leveraging internal reinforcement signals, RLVR mitigates catastrophic forgetting and often achieves superior performance. Nevertheless, its performance can degrade when tasks extend beyond the initial model’s knowledge base. We evaluate these two paradigms across a diverse set of vision-and-language tasks, with our key findings summarized in Figure 1. The results indicate that SFT is more effective for tasks that fall outside the pre-training distribution of LVLMs, whereas RLVR yields superior performance on tasks that align with its pre-existing knowledge base. We also provide a case study in Section 3.2 to illustrate this phnomenon. Although a sequential SFT → RLVR pipeline tries to leverage their complementary strengths, this two-stage method incurs the combined computational cost of both stages and remains susceptible to catastrophic forgetting during the initial SFT phase.
To address the aforementioned limitations, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified, single-stage post-training paradigm designed to integrate the complementary advantages of SFT and RLVR. We begin with an analysis of the underlying objectives and gradients of SFT and RLVR. We theoretically demonstrate that their gradients share similar patterns, enabling them to be integrated into a single, unified objective, the ViSurf objective. Unlike prior methods [zhang2025policy, ma2025learning] that combine SFT and RLVR through simple objective addition, ViSurf offers a unified perspective for these two training paradigms. While subtle differences exist, the gradient of ViSurf objective can be interpreted as a composite of the gradients from both SFT and RLVR. Based on the theoretical analysis, we design three reward control strategies to stabilize training. Specifically, for the ground-truth labels, we i) aligning them with rollouts preference, ii) eliminating thinking reward for them, and iii) smoothing the reward for them. The implementation of ViSurf is consequently simplified to interleaving ground-truth demonstrations with on-policy rollouts in a unified training phase.
@@
We conduct extensive experiments across various domains, with a comparative summary presented in Figure 2. The results indicate that our method, ViSurf, achieves superior performance than SFT, RLVR and SFT → RLVR. Models trained with ViSurf also exhibit reasoning abilities similar to those established in previous works [liu2025seg, liu2025visionreasoner]. Furthermore, ViSurf successfully mitigates catastrophic forgetting, as evidenced by its stable performance on VQA tasks. The ablation study confirms the critical contribution of the proposed reward control mechanism. Additional in-depth analysis provides empirical validation for our theoretical analysis and offers insights into the operational principles of ViSurf. Our contributions are summarized as follows:
  • Based on our theoretical analysis, we propose ViSurf, a unified post-training paradigm that leverages the complementary benefits of SFT and RLVR.
  • We design three reward control strategies to stabilize the training process, the necessity of which is validated through ablation studies.
  • Empirically, ViSurf outperforms existing post-training methods (SFT, RLVR, and SFT → RLVR). Additional analyses offer a thorough understanding of its mechanics.
@@ 2.1 Supervised Fine-tuning for LVLMs
Supervised Fine-Tuning (SFT) is a predominant paradigm for training Large Vision-Language Models (LVLMs). This approach involves fine-tuning pre-trained models on expert-annotated data. The pioneering work in this area is LLaVA [liu2023visual], which has inspired numerous subsequent models. Notable examples include the LLaVA-series [li2024llavaov, liu2024improved], QwenVL-series [wang2024qwen2vl, bai2025qwen25vl], MGM-series [li2024mini, zhong2024lyra, wang2025mgm], and InternVL [chen2024internvl], all of which adopt this paradigm. SFT has proven particularly effective for adapting LVLMs to diverse downstream applications, such as image quality assessment [you2025teaching], visual counting [deitke2024molmo], and autonomous driving [xu2025drivegpt4].
@@ 2.2 Reinforcement Learning for LVLMs
Reinforcement Learning (RL) is a standard method for fine-tuning Large Vision-Language Models (LVLMs). Among RL algorithms, Direct Preference Optimization (DPO) [rafailov2023direct] relies on pre-collected human preference datasets, which can be costly to produce. Similarly, Proximal Policy Optimization (PPO) [schulman2017proximal] requires a well-trained reward model to evaluate responses generated by the policy. Recently, RLVR algorithms, such as GRPO [shao2024deepseekmath] and DAPO [yu2025dapo], have gained attention for their ability to assess model outputs against objective criteria. This approach reduces the dependency on manually annotated data and pre-trained reward models. The effectiveness of RLVR for LVLMs has been demonstrated in recent works [liu2025seg, liu2025visual, liu2025visionreasoner, liu2025empowering, huang2025vision] such as SegZero [liu2025seg] and VisualRFT [liu2025visual].
@@ 3 ViSurf
We begin by analyzing the objective functions of SFT and RLVR in Section 3.1. A case study in Section 3.2 then highlights the limitations of both methods. To address these limitations, we introduce ViSurf, detailing its design and relationship to SFT and RLVR in Section 3.3. Subsequently, Section 3.4 presents three novel mechanisms for reward control during training, and Section 3.5 provides an analysis of the optimization process.
@@ 3.1 Preliminary
$\mathcal{L}_{\text{SFT}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\text{input}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}\left[\log\pi_{\theta}(y\mid v,t)\right],$ (1)
where $y$ corresponds to $(v,t)$. A more precise notation would be $(v,t,{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y})\sim\text{zip}(\mathcal{D}_{\text{input}},{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\mathcal{D}_{\text{label}}})$. Nevertheless, we retain the current notation, $(v,t)\sim\mathcal{D}_{\text{input}},{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}$, for clarity and ease of comparison in the subsequent discussion.
Reinforcement Learning with Verifiable Rewards (RLVR). We illustrate RLVR using the on-policy Group Relative Policy Optimization (GRPO) algorithm [shao2024deepseekmath]. GRPO optimizes the policy $\pi_{\theta}$ using a verifiable reward function, which typically combines measures of output format and accuracy [guo2025deepseek, liu2025visionreasoner, liu2025seg]. For a given input $(v_{i},t_{i})\in\mathcal{D}_{\text{input}}$, the old policy $\pi_{\theta_{old}}$ (from a previous optimization step) generates a group of $G$ rollouts $\{o_{j}\}^{G}_{j=1}$ by sampling with different random seeds. Each rollout $o_{j}$ is then evaluated by a reward function $r(\cdot)$, resulting in a set of rewards $\{r(o_{j})\}^{G}_{j=1}$. The advantage for each rollout is subsequently computed as follows:
@@
$\mathcal{L}_{\mathrm{RLVR}}(\theta)$ $=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}\left[\frac{1}{G}\sum_{j=1}^{G}\min\Bigg\{\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)}\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}},\right$. (3)
$\hskip 17.00024pt\hskip 17.00024pt\left.\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)},\;1-\epsilon,\;1+\epsilon\Biggr)\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\Bigg\}\right],$
where the $\epsilon$ is a constant that controls the clipping boundary. For simplicity, both in equation and in our practical implementation, we omit the KL divergence term.
@@ 3.2 Case Study: Non-Object Scenarios
We present a case study on non-object referring expression segmentation. In this task, instructions comprise both correct expressions (referring to existing objects) and incorrect ones (where the referred objects are absent). For this study, we utilize the VisionReasoner model [liu2025visionreasoner], which is initialized with Qwen2.5VL [bai2025qwen25vl] and SAM2 [ravi2024sam2] and has demonstrated strong performance on standard referring segmentation tasks. Figure 3(a) illustrates the non-object segmentation and the architecture of VisionReasoner. Detailed experimental settings are demonstrated Section 4.1.
@@
The gRefCOCO [liu2023gres] serves as the benchmark for this scenario. We compare gIoU (average IoU across images) and N-Acc (accuracy in identifying non-existent object). Quantitative and qualitative results are presented in Figure 3(b)-(c). Our analysis reveals that model trained by SFT achieves suboptimal gIoU performance but tends to learn to correctly identify the absence of objects. In contrast, model trained with RLVR exhibits a more coherent reasoning process and attain higher overall gIoU scores, yet they often generate object masks even when no relevant object is present. This limitation arises because the RLVR model, relying solely on self-rollouts, lacks the corrective mechanism necessary to produce a “no object” output. In essence, RLVR enhances overall performance through internal self-guidance, whereas SFT provides crucial external guidance when self-rollouts fail. This analysis leads to a key question: how can we integrate the benefits of both training paradigms in a single training stage efficiently?
@@ 3.3 Combining SFT and RLVR
$\footnotesize\nabla_{\theta}\mathcal{L}_{\mathrm{SFT}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\mathrm{label}}}\end{subarray}}\left[\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\right].$ (4)
The gradient of RLVR can be derived from Equation 3 using approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ and log-derivative trick. We also omit clip operation for simplicity:
$\nabla_{\theta}\mathcal{L}_{\mathrm{RLVR}}(\theta)$ $=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}$ (5)
$\qquad\left[\frac{1}{G}\sum_{j=1}^{G}{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)\right]_{\,\theta\approx\theta_{\mathrm{old}}}.$
@@
$\mathcal{L}_{\mathrm{{ViSurf}}}$ $(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}$ (8)
$\Bigg[\frac{1}{G+1}\Bigg(\sum_{j=1}^{G}\min\Bigg\{\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)}\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}},$
$\qquad\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)},1-\epsilon,1+\epsilon\Biggr)\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\Bigg\}$
$+\min\Bigg\{\frac{\pi_{\theta}(y\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(y\mid v,t)}\,{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\hat{A}_{y}},$
$\qquad\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(y\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(y\mid v,t)},1-\epsilon,1+\epsilon\Biggr)\,{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\hat{A}_{y}}\Bigg\}\Bigg)\Bigg].$
@@
Input: policy model $\pi_{\theta}$; reward function $\text{r}(\cdot)$; input data $\mathcal{D}_{\mathrm{input}}$; label data $\mathcal{D}_{\mathrm{label}}$
@@
  Sample a mini-batch $\mathcal{B}_{\mathrm{input}}$ and corresponding $\mathcal{B}_{\mathrm{label}}$;
  Update the old policy model $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$;
  Sample $G$ outputs $\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\text{old}}}()$ for each $(v,t)\in\mathcal{B}_{\mathrm{input}}$;
  Compute rewards $\{\text{r}(o_{j})\}_{j=1}^{G}$ for each sampled output $o_{i}$;
  Compute rewards $\text{r}(y)$ for label $y\in\mathcal{B}_{\mathrm{label}}$;
  Compute $\hat{A}_{j}$ and $\hat{A}_{y}$ through relative advantage estimation;
  Update the policy model $\pi_{\theta}$ using Equation 8;

### v3

@@ Abstract
Post-training Large Vision-and-Language Models (LVLMs) typically involves Supervised Fine-Tuning (SFT) for knowledge injection or Reinforcement Learning with Verifiable Rewards (RLVR) for performance enhancement. However, SFT often leads to sub-optimal performance, while RLVR remains constrained by the model’s internal knowledge base. While a sequential SFT $\rightarrow$ RLVR pipeline can be used, it introduces significant computational overhead and suffers from catastrophic forgetting. To address these limitations, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified, single-stage paradigm that integrates the strengths of both SFT and RLVR. By analyzing their training objectives, we establish a unified framework that injects ground-truth labels directly into RLVR rollouts, facilitating simultaneous external supervision and internal reinforcement. Furthermore, we introduce three novel reward control strategies to ensure training stability and optimization. Extensive experiments demonstrate that ViSurf consistently outperforms standalone SFT, RLVR, and the traditional two-stage pipeline across diverse benchmarks. In-depth analysis corroborates these findings, validating the derivation and design principles of ViSurf. Our code, data and models will be publicly available.
Machine Learning
@@ 1 Introduction
Developing Large Vision-and-Language Models (LVLMs) capable of excelling across diverse visual perception tasks represents a pivotal frontier in visual intelligence. To date, research has predominantly relied on two distinct post-training paradigms: Supervised Fine-Tuning (SFT) (Liu et al., 2023b; Wang et al., 2024c; Bai et al., 2025) and Reinforcement Learning with Verifiable Rewards (RLVR) (Liu et al., 2025b, c).
However, these paradigms exhibit a fundamental dichotomy in their strengths and weaknesses. SFT directly optimizes models using expert-annotated data, providing explicit external guidance that enables the model to approximate target distributions. Despite this, SFT often yields sub-optimal generalization and is prone to catastrophic forgetting of pre-trained knowledge. Conversely, RLVR, particularly on-policy methods like Group Relative Policy Optimization (GRPO) (Shao et al., 2024) and Dynamic Sampling Policy Optimization (DAPO) (Yu et al., 2025), leverages internal reinforcement signals to optimize policies based on rollout evaluations. While RLVR effectively mitigates forgetting and enhances performance, its efficacy is fundamentally capped by the model’s inherent knowledge boundaries; performance typically stagnates when tasks extend beyond the initial policy’s distribution. Our evaluation across diverse vision-language benchmarks, summarized in Figure 1, confirms this phenomenon: SFT excels in out-of-distribution knowledge acquisition, whereas RLVR thrives on tasks aligning with pre-existing capabilities (see Section 3.2 for a detailed case study). While a sequential SFT $\rightarrow$ RLVR pipeline attempts to marry these strengths, it introduces prohibitive computational overhead and remains vulnerable to forgetting during the initial SFT phase.
To overcome these challenges, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified, single-stage paradigm that integrates the complementary advantages of SFT and RLVR. We provide a rigorous analysis of the underlying objectives and gradients for both methods, theoretically demonstrating that their shared gradient patterns allow for integration into a singular ViSurf objective. Unlike existing approaches (Zhang et al., 2025; Ma et al., 2025) that simply sum the SFT and RLVR losses, ViSurf offers a theoretically grounded, unified perspective. While subtle differences exist, the gradient of ViSurf objective can be interpreted as a composite of the gradients from both SFT and RLVR. To ensure training stability within this unified framework, we introduce three novel reward control strategies for ground-truth labels: (i) preference alignment with policy rollouts, (ii) exclusion of ”thinking” rewards for static labels, and (iii) reward smoothing to prevent optimization spikes. Consequently, the implementation of ViSurf is streamlined into an elegant process: interleaving ground-truth demonstrations with on-policy rollouts within a single, unified training phase.
@@
We evaluate ViSurf across diverse domains, with a comparative summary presented in Figure 2. The results demonstrate that ViSurf consistently outperforms SFT, RLVR, and the SFT $\rightarrow$ RLVR pipeline. Notably, our method mitigates catastrophic forgetting, as evidenced by its stable performance on VQA benchmarks. Furthermore, our ablation study confirms the critical role of the proposed reward control mechanism, while in-depth analysis provides empirical support for our theoretical framework and offers insights into the operational principles of ViSurf. Our contributions are summarized as follows:
  • Unified Framework: We introduce ViSurf, a theoretically grounded, single-stage paradigm that unifies SFT and RLVR to achieve simultaneous knowledge injection and internal reinforcement.
  • Stability Optimization: We design three novel reward control strategies to stabilize joint optimization, effectively balancing ground-truth guidance with on-policy exploration.
  • State-of-the-Art Performance: ViSurf consistently outperforms standalone SFT, RLVR, and two-stage pipelines across diverse benchmarks, supported by in-depth analysis of its synergistic mechanics.
@@ 2.1 Supervised Fine-tuning for LVLMs
Supervised Fine-Tuning (SFT) has emerged as the predominant paradigm for advancing Large Vision-Language Models (LVLMs), typically involving the refinement of pre-trained models using expert-annotated datasets. Prominent examples include the LLaVA-series (Liu et al., 2023b; Li et al., 2024a; Liu et al., 2024), QwenVL-series (Wang et al., 2024c; Bai et al., 2025), MGM-series (Li et al., 2024b; Zhong et al., 2024; Wang et al., 2025), and InternVL (Chen et al., 2024b), all of which leverage the SFT methodology. Beyond foundational training, SFT has demonstrated remarkable efficacy in tailoring LVLMs to specialized downstream tasks, such as image quality assessment (You et al., 2025) and autonomous driving (Xu et al., 2025).
@@ 2.2 Reinforcement Learning for LVLMs
Beyond supervised methods, Reinforcement Learning (RL) has become a cornerstone for aligning Large Vision-Language Models (LVLMs) with human expectations. Established techniques like DPO (Rafailov et al., 2023) necessitate extensive human-labeled preference datasets, which are often prohibitively expensive to curate. Similarly, PPO (Schulman et al., 2017) relies on a robust reward model to supervise the policy’s outputs. To mitigate these constraints, Reinforcement Learning from Verifiable Rewards (RLVR) algorithms, such as GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025), leverage verification functions to evaluate model performance. The potency of this paradigm for LVLMs is underscored by recent advancements (Liu et al., 2025b, d, c, a; Huang et al., 2025), most notably in SegZero (Liu et al., 2025b) and VisualRFT (Liu et al., 2025d).
@@ 3 ViSurf
We begin by analyzing the objective functions of SFT and RLVR in Section 3.1. A case study in Section 3.2 then highlights the limitations of both methods. To address these limitations, we introduce ViSurf, detailing its design and its theoretical alignment with SFT and RLVR in Section 3.3. Subsequently, Section 3.4 presents three novel mechanisms for reward control during training, and Section 3.5 a rigorous analysis of the underlying optimization dynamics.
@@ 3.1 Preliminary
$\mathcal{L}_{\text{SFT}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\text{input}}\\ {\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}\left[\log\pi_{\theta}(y\mid v,t)\right],$ (1)
where $y$ corresponds to $(v,t)$. A more precise notation would be $(v,t,{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}y})\sim\text{zip}(\mathcal{D}_{\text{input}},{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}\mathcal{D}_{\text{label}}})$. Nevertheless, we retain the current notation, $(v,t)\sim\mathcal{D}_{\text{input}},{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}y\sim\mathcal{D}_{\text{label}}}$, for clarity and ease of comparison in the subsequent discussion.
Reinforcement Learning with Verifiable Rewards (RLVR). We illustrate RLVR using the on-policy Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024). GRPO optimizes the policy $\pi_{\theta}$ using a verifiable reward function, which typically combines measures of output format and accuracy (Guo et al., 2025; Liu et al., 2025c, b). For a given input $(v_{i},t_{i})\in\mathcal{D}_{\text{input}}$, the old policy $\pi_{\theta_{old}}$ (from a previous optimization step) generates a group of $G$ rollouts $\{o_{j}\}^{G}_{j=1}$ by sampling with different random seeds. Each rollout $o_{j}$ is then evaluated by a reward function $r(\cdot)$, resulting in a set of rewards $\{r(o_{j})\}^{G}_{j=1}$. The advantage for each rollout is subsequently computed as follows:
@@
$\mathcal{L}_{\mathrm{RLVR}}(\theta)$ $=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}\left[\frac{1}{G}\sum_{j=1}^{G}\min\Bigg\{\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)}\,{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\hat{A}_{j}},\right$. (3)
$\hskip 17.00024pt\hskip 17.00024pt\left.\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)},\;1-\epsilon,\;1+\epsilon\Biggr)\,{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\hat{A}_{j}}\Bigg\}\right],$
where the $\epsilon$ is a constant that controls the clipping boundary. For simplicity, both in equation and in our practical implementation, we remove the KL divergence term.
@@ 3.2 Case Study: Non-Object Scenarios
We conduct a case study on non-object referring expression segmentation, a challenging task where instructions include both valid expressions and “incorrect” ones referring to non-existent objects. For this study, we utilize the VisionReasoner architecture (Liu et al., 2025c), which is initialized with Qwen2.5-VL (Bai et al., 2025) and SAM2 (Ravi et al., 2024) and has demonstrated strong performance on standard referring segmentation tasks. Figure 3(a) illustrates the non-object segmentation and the architecture of VisionReasoner. Experimental configurations are detailed in Section 4.1.
@@
The gRefCOCO (Liu et al., 2023a) serves as the benchmark for this scenario. We compare gIoU (average IoU across images) and N-Acc (accuracy in identifying non-existent object). Quantitative and qualitative results are presented in Figure 3(b)-(c). Our analysis reveals that model trained by SFT achieves suboptimal gIoU performance but tends to learn to correctly identify the absence of objects. In contrast, model trained with RLVR attain higher overall gIoU scores, yet they often generate object masks even when no relevant object is present. This limitation arises because the RLVR model, relying solely on self-rollouts, lacks the corrective mechanism necessary to produce a correct “no object” output. In essence, RLVR drives performance gains through self-exploration, whereas SFT provides the external grounding necessary when exploration fails. This dichotomy motivates our central research question: how can we efficiently synthesize the strengths of both training paradigms within a unified training stage?
@@ 3.3 Combining SFT and RLVR
$\footnotesize\nabla_{\theta}\mathcal{L}_{\mathrm{SFT}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}y\sim\mathcal{D}_{\mathrm{label}}}\end{subarray}}\left[\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\right].$ (4)
In our practical implementation, we adopt a single-sampling, single-updating paradigm, which effectively bypasses the need for clipping operations. Therefore, the gradient of RLVR can be derived from Equation 3 using approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ and log-derivative trick:
$\nabla_{\theta}\mathcal{L}_{\mathrm{RLVR}}(\theta)$ $=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}$ (5)
$\qquad\left[\frac{1}{G}\sum_{j=1}^{G}{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)\right]_{\,\theta\approx\theta_{\mathrm{old}}}.$
@@
$\mathcal{L}_{\mathrm{{ViSurf}}}$ $(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\\ {\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}$ (8)
$\Bigg[\frac{1}{G+1}\Bigg(\sum_{j=1}^{G}\min\Bigg\{\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)}\,{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\hat{A}_{j}},$
$\qquad\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)},1-\epsilon,1+\epsilon\Biggr)\,{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\hat{A}_{j}}\Bigg\}$
$+\min\Bigg\{\frac{\pi_{\theta}(y\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(y\mid v,t)}\,{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}\hat{A}_{y}},$
$\qquad\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(y\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(y\mid v,t)},1-\epsilon,1+\epsilon\Biggr)\,{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}\hat{A}_{y}}\Bigg\}\Bigg)\Bigg].$
@@
Input: policy model $\pi_{\theta}$; reward function $\text{r}(\cdot)$;
input data $\mathcal{D}_{\mathrm{input}}$; label data $\mathcal{D}_{\mathrm{label}}$
@@
  Sample a mini-batch $\mathcal{B}_{\mathrm{input}}$ and corresponding $\mathcal{B}_{\mathrm{label}}$
 Update the old policy model $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
 Sample $G$ outputs $\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\text{old}}}()$
    for each $(v,t)\in\mathcal{B}_{\mathrm{input}}$
 Compute rewards $\{\text{r}(o_{j})\}_{j=1}^{G}$ for each sampled output $o_{j}$
 Compute rewards $\text{r}(y)$ for label $y\in\mathcal{B}_{\mathrm{label}}$
 Compute $\hat{A}_{j}$ and $\hat{A}_{y}$ through relative advantage estimation
 Update the policy model $\pi_{\theta}$ using Equation 8